# Using Dynamic Safety Margins as Control Barrier Functions

### v4

@@ Abstract
{IEEEkeywords}
Constrained control, Control barrier functions, Predictive control for nonlinear systems.
1 Introduction
[…]
@@ 1 Introduction
Figure 1: Comparison of CBF-based control (top) with RG-based control (bottom). The CBF approach filters a nominal input signal $\boldsymbol{\mathbf{u}}_{n}$ to obtain a safe input $\boldsymbol{\mathbf{u}}$. The RG approach a nominal reference signal $\boldsymbol{\mathbf{r}}_{n}$ to obtain a safe virtual reference $\boldsymbol{\mathbf{r}}$ for the nominal controller.
1.1 Related Works
@@ 1.2 Contributions
  • We leverage the control-sharing property presented in [9] to define a natural generalization of control barrier functions (CBFs) to (vector-valued) CBFs.
  • We show that dynamic safety margins (DSMs) are (vector-valued) CBFs for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM. Additionally, we show that the relative degree of the constraints is irrelevant for DSMs by construction.
@@
  • We extend the Lyapunov-based DSMs presented in [24] to inadmissible references. This extension is needed to use them as (vector-valued) CBFs. Moreover, we show that any class $\mathscr{K}$ function satisfies the CBF condition of Lyapunov-based DSMs.
@@
The rest of the paper is organized as follows: Section 2 provides formal definitions for control invariance, CBFs, and DSMs; Section 3 proves that DSMs are CBFs, shows the relative degree of the constructed CBF is irrelevant for safety, formulates the CBF-based optimization problem, which is a quadratic program (QP) for polyhedral input constraints, and details some conditions that guarantee the DSM-CBF policy is locally Lipschitz continuous; Section 4 provides a systematic approach based on Lyapunov functions to construct DSMs and highlights some desirable properties that such DSMs have; Section 5 summarizes the presented CBF synthesis approach and contextualizes it with respect to the existing reference governor literature; Section 6 presents the numerical simulations and comparisons; Section 7 concludes the paper.
1.3 Notation
Throughout the paper, only vector quantities are presented in bold font: $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$. The notation $\boldsymbol{\mathbf{x}};~\boldsymbol{\mathbf{y}}]$ denotes the vertical stacking of two vectors $\boldsymbol{\mathbf{x}}$ and $\boldsymbol{\mathbf{y}}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}$ if it is strictly increasing and $\alpha(0)=0$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}=\text{closure}(\mathcal{S})\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Given a subset $\mathcal{C}\subset\mathbb{R}^{n}$ of the state space $\mathbb{R}^{n}$, we generally denote by $\tilde{\mathcal{C}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ its counterpart in the augmented state-reference space. Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ is a regular point of $h$ if $\partial h(\boldsymbol{\mathbf{x}})/\partial\boldsymbol{\mathbf{x}}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\boldsymbol{\mathbf{x}}\in\mathcal{D}$ such that $h(\boldsymbol{\mathbf{x}})=c$ is a regular point.
2 Preliminaries
@@
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}},$ (1)
where $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions. The system is subject to state $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and input $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ constraints, where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{U}\subset\mathbb{R}^{m}$ are closed sets.
2.1 Control Invariance
@@
A set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if, $\forall\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}$, there exists an admissible input signal $\boldsymbol{\mathbf{u}}:[0,\infty)\to\mathcal{U}$ such that at least one of the trajectories of the closed-loop system (1), with $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$, satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t$ in its interval of existence $\mathcal{J}(\boldsymbol{\mathbf{x}}_{0})\subset[0,\infty)$.
Note that this property is sometimes called viability [28] or weak control invariance [27] because it doesn’t require existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s theorem characterizing control invariance in terms of Bouligand’s tangent cone.
@@
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\exists\boldsymbol{\mathbf{u}}\in\mathcal{U},\quad f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\in\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}}),$ (2)
where $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is the tangent cone to $\mathcal{C}$ in $\boldsymbol{\mathbf{x}}$ [27, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\boldsymbol{\mathbf{x}})$ is cumbersome. Fortunately, when the set $\mathcal{C}$ is the 0-superlevel set of a continuously differentiable function, the tangent cone on the boundary is reduced to the tangent half-space shifted to the origin and we have the following characterization.
Theorem 2 ([30, Proposition 4.3.7])
Let $h:\mathcal{D}\to\mathbb{R}$ be a continuously differentiable function with regular value 0. The set $\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}:h(\boldsymbol{\mathbf{x}})\geq 0\}$ is control invariant if and only if
$\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C},\quad\exists\boldsymbol{\mathbf{u}}\in\mathcal{U},\quad L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0,$ (3)
where $L_{f}h$ and $L_{g}h$ are the Lie derivatives of $h$ along $f$ and $g$, respectively.
Note that [30, Proposition 4.3.7] is more general than Theorem 2: it allows the set $\mathcal{C}$ to be defined using multiple continuously differentiable functions. We use this fact in upcoming results.
@@
2.2 Control Barrier Functions
Control barrier functions certify the control invariance of their superlevel set.
Definition 2 ([6])
@@ 2.2 Control Barrier Functions
A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ is a control barrier function (CBF) if there exists a class $\mathscr{K}$ function $\alpha:[0,\infty)\to[0,\infty)$ such that
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right]\geq-\alpha\big(h(\boldsymbol{\mathbf{x}})\big),$ (4)
where
$\mathcal{C}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}.$ (5)
@@
With this definition, the conditions of Theorem 2 were extended to the entire set $\mathcal{C}$, as shown in the next result.
Lemma 1
Assume $\mathcal{U}\!\subset\!\mathbb{R}^{m}$ is compact. Let $\mathcal{C}\!\subset\!\mathbb{R}^{n}$, defined as in (5), be compact and let 0 be a regular value of $h$. The set $\mathcal{C}$ is control invariant if and only if $h$ is a CBF.
Proof 2.3.
Although the proof is based on [6], they showed this result for barrier functions, not control barrier functions. Let $\mathcal{C}\subset\mathbb{R}^{n}$ be control invariant and define the set-valued map $C(c)=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid 0\leq h(\boldsymbol{\mathbf{x}})\leq c\}$. Note that for any $c\geq 0$, $C(c)\subset\mathcal{C}$ is compact in $\mathbb{R}^{n}$. Next, define the function $\hat{\alpha}:[0,\infty)\to\mathbb{R}$
$\hat{\alpha}(c)=-\inf_{\boldsymbol{\mathbf{x}}\in C(c)}\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right].$ (6)
Since $\mathcal{U}$ is compact, it follows by the extreme value theorem that $\xi:\mathcal{D}\to\mathbb{R}$, defined as
$\xi(\boldsymbol{\mathbf{x}})=\sup_{\boldsymbol{\mathbf{u}}\in\mathcal{U}}\left[L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\right],$ (7)
is continuous on $\mathcal{D}$. Moreover, it can be shown that the set-valued map $C$ is both uppper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [31], that the value function $\hat{\alpha}$ is continuous on $0,\infty)$. By construction, $\hat{\alpha}$ is non-decreasing. By Theorem 2, for any $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq 0$. Thus, $\forall\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}=\mathcal{C}(0)$, $\xi(\boldsymbol{\mathbf{x}})\geq 0$. As a result,
$\hat{\alpha}(0)=-\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}}\xi(\boldsymbol{\mathbf{x}})\leq 0.$ (8)
There always exists a class $\mathscr{K}$ function $\alpha:[0,\infty)\to[0,\infty)$ that upper-bounds $\hat{\alpha}$. Finally, let $\boldsymbol{\mathbf{x}}\in\mathcal{C}$ be given and note
$\xi(\boldsymbol{\mathbf{x}})\geq\inf_{\boldsymbol{\mathbf{x}}_{0}\in\mathcal{C}(h(\boldsymbol{\mathbf{x}}))}\xi(\boldsymbol{\mathbf{x}}_{0})=-\hat{\alpha}\big(h(\boldsymbol{\mathbf{x}})\big)\geq-\alpha\big(h(\boldsymbol{\mathbf{x}})\big).$
So, we conclude that $h$ is a CBF.
On the other hand, assume $h$ is a CBF, let $\alpha\in\mathscr{K}$ be such that (4) holds, and let $\boldsymbol{\mathbf{x}}\in\partial\mathcal{C}$ be given. By compactness of $\mathcal{U}$, there exists $\boldsymbol{\mathbf{u}}\in\mathcal{U}$ such that $L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\geq-\alpha\big(h(\boldsymbol{\mathbf{x}})\big)=0$, where the last equality holds because $h(\boldsymbol{\mathbf{x}})=0$. Since $\mathcal{C}$ is closed and 0 is a regular value of $h$, we can apply Theorem 2 to conclude $\mathcal{C}$ is control invariant.
Notice compactness of $\mathcal{U}$ implies $\xi:\mathcal{D}\to\mathbb{R}$ is continuous, and therefore $\hat{\alpha}:[0,\infty)\to\mathbb{R}$. Without it, we lose necessity in Lemma 1. However, the sufficiency in Lemma 1 (i.e., the CBF certifies control invariance of its compact superlevel set) is maintained even when the compactness of $\mathcal{U}$ is exchanged for the less restrictive condition
$\forall\boldsymbol{\mathbf{x}}\in\mathcal{C},\quad\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})\neq\emptyset,$ (9)
where
$\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})\!=\!\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h(\boldsymbol{\mathbf{x}})+L_{g}h(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big(h(\boldsymbol{\mathbf{x}})\big)\!\geq\!0\}.$ (10)
In this work we drop the compactness requirement on $\mathcal{U}$ because only the sufficiency statement in Lemma 1 is relevant and we will design CBFs for which (9) holds. Compactness of $\mathcal{C}$ is required for the necessity (control invariant $\mathcal{C}$ implies $h$ is a CBF) of Lemma 1, but not for the sufficiency ($h$ is a CBF implies $\mathcal{C}$ is control invariant) of Lemma 1.
@@
CBFs can be used to design add-on modules that enforce constraint satisfaction by filtering the control input. Specifically, let $h$ be a CBF such that $\mathcal{C}\subset\mathcal{X}$ and consider the closed-loop system
$\dot{\boldsymbol{\mathbf{x}}}=f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\beta(\boldsymbol{\mathbf{x}}),$ (11)
@@
$\beta(\boldsymbol{\mathbf{x}})=\underset{\boldsymbol{\mathbf{u}}\in\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})}{\operatorname{argmin}}\>\|\boldsymbol{\mathbf{u}}-\kappa(\boldsymbol{\mathbf{x}})\|^{2},$ (12)
and $\kappa:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a nominal control policy with desirable closed-loop performance. When (9) holds, the optimization problem (12) is always feasible, making $\beta$ well-defined. If, in addition, the policy $\beta$ is locally Lipschitz continuous, there exists a unique solution $\boldsymbol{\mathbf{x}}(t)$ to the closed-loop system (11), and it satisfies $\boldsymbol{\mathbf{x}}(t)\!\in\!\mathcal{C}\!\subset\!\mathcal{X}$. In upcoming sections, we provide conditions under which $\beta$ is locally Lipschitz continuous.
Often, the sets $\mathcal{X}$ and $\mathcal{U}$ are the intersection of multiple state and input constraints. In these cases, it is usually easier to design a separate CBF for each constraint and then enforce all the conditions in one CBF-based program. Unfortunately, there is typically no guarantee that feasibility is retained when all the constraints are superimposed. The following property certifies feasibility is retained in these cases
Definition 2.4 ([9, Definition 2]).
Consider $n_{c}\in\mathbb{N}$ continuously differentiable functions $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and let $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}\subset\mathbb{R}^{n}$, with $\mathcal{C}=\cap_{i=1}^{n_{c}}\mathcal{C}_{i}$. The functions $h_{i}$ are said to have the control-sharing property if there exists a class $\mathscr{K}$ function $\alpha$ satisfying $\forall\boldsymbol{\mathbf{x}}\in\mathcal{C}$, $\exists\boldsymbol{\mathbf{u}}\in\mathcal{U}$, such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[L_{f}h_{i}(\boldsymbol{\mathbf{x}})+L_{g}h_{i}(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha\big(h_{i}(\boldsymbol{\mathbf{x}})\big)\right]\!\geq\!0.$ (13)
In this paper, we call a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ whose channels $h_{i}$ have the control-sharing property a (vector-valued) control barrier function.
The functions $h_{i}$ in Definition 2.4 are deliberately not labeled as CBFs individually (as opposed to [9, Definition 2]) because such a requirement is unnecessarily limiting. In fact, if 0 is a regular value of each $h_{i}$, Lemma 1 also holds for the set $\mathcal{C}$ defined as the intersection of the $\mathcal{C}_{i}$ even if each $\mathcal{C}_{i}$ is individually not compact. The optimization-based policy $\beta$ can also be used with (vector-valued) CBFs with the following admissible input set

### v5

@@ Abstract
I Introduction
@@
Figure 1: Comparison of CBF-based control (top) with RG-based control (bottom). The CBF approach filters a nominal input signal $\bm{u}_{n}$ to obtain a safe input $\bm{u}$. The RG approach a nominal reference signal $\bm{r}_{n}$ to obtain a safe virtual reference $\bm{r}$ for the nominal controller.
I-A Related Works
@@ I-B Contributions
  • We leverage the control-sharing property presented in [44] to define a natural generalization of traditional real-valued control barrier functions (CBFs) to vector-valued CBFs.
  • We show that dynamic safety margins (DSMs) are CBFs for an augmented system consisting of the concatenation of the state vector and the reference of the prestabilizing controller used to construct the DSM. Additionally, we show that the relative degree of the constraints is irrelevant for DSMs by construction.
@@
  • We extend the Lyapunov-based DSMs presented in [38] to inadmissible references, needed to use them as CBFs. Moreover, we show that any class $\mathscr{K}$ function satisfies the CBF condition of Lyapunov-based DSMs.
@@
The rest of the paper is organized as follows: Section II provides formal definitions for control invariance, CBFs, and DSMs; Section III proves that DSMs are CBFs, shows the relative degree of the constructed CBF is irrelevant for safety, formulates the CBF-based optimization problem, which is a quadratic program (QP) for polyhedral input constraints, and details some conditions that guarantee the DSM-CBF policy is locally Lipschitz continuous; Section IV provides a systematic approach based on Lyapunov functions to construct DSMs and highlights some desirable properties that such DSMs have; Section V summarizes the presented CBF synthesis approach and contextualizes it with respect to the existing reference governor literature; Section VI presents the numerical simulations and comparisons; Section VII concludes the paper.
I-C Notation
Throughout the paper, only vector quantities are presented in bold font: $\bm{x}\in\mathbb{R}^{n}$. The notation $\bm{x};~\bm{y}]$ denotes the vertical stacking of two vectors $\bm{x}$ and $\bm{y}$ into a column vector. A continuous function $\alpha:[0,\infty)\to[0,\infty)$ is said to belong to class $\mathscr{K}$ if it is strictly increasing and $\alpha(0)=0$. The symbol $\subset$ denotes set inclusion, but not necessarily strict inclusion. Given a set $\mathcal{S}$, we denote by $\operatorname{Int}\left(\mathcal{S}\right)$ the interior of $\mathcal{S}$ and by $\partial\mathcal{S}$ the boundary of $\mathcal{S}$ (i.e. $\partial\mathcal{S}=\text{closure}(\mathcal{S})\setminus\operatorname{Int}\left(\mathcal{S}\right)$). Given a subset $\mathcal{C}\subset\mathbb{R}^{n}$ of the state space $\mathbb{R}^{n}$, we generally denote by $\tilde{\mathcal{C}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ its counterpart in the augmented state-reference space. Let $h:\mathcal{D}\to\mathbb{R}$ be continuously differentiable. We say $\bm{x}\in\mathcal{D}$ is a regular point of $h$ if $\partial h(\bm{x})/\partial\bm{x}\neq 0$. We say $c\in\mathbb{R}$ is a regular value of $h$ if every $\bm{x}\in\mathcal{D}$ such that $h(\bm{x})=c$ is a regular point.
II Preliminaries
@@
$\dot{\bm{x}}=f(\bm{x})+g(\bm{x})\bm{u},$ (1)
where $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions. The system is subject to state $\bm{x}\in\mathcal{X}$ and input $\bm{u}\in\mathcal{U}$ constraints, where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{U}\subset\mathbb{R}^{m}$ are closed sets.
II-A Control Invariance
@@
A set $\mathcal{C}\subset\mathbb{R}^{n}$ is control invariant if, $\forall\bm{x}_{0}\in\mathcal{C}$, there exists an admissible, essentially bounded input signal $u:[0,\infty)\to\mathcal{U}$ such that at least one of the trajectories of the closed-loop system (1), with $x(0)=\bm{x}_{0}$, satisfies $x(t)\in\mathcal{C}$ for all $t$ in its interval of existence $\mathcal{J}(\bm{x}_{0})\subset[0,\infty)$.
Note that this property is sometimes called weak control invariance [9] or viability [6] and that it doesn’t require existence and uniqueness of the trajectory. Nonetheless, it will simplify the presentation of the upcoming results. Assumptions for the existence and uniqueness of solutions will be presented separately. Central to the upcoming discussion is Nagumo’s theorem characterizing control invariance in terms of Bouligand’s tangent cone.
@@
$\forall\bm{x}\in\mathcal{C},\quad\exists\bm{u}\in\mathcal{U},\quad f(\bm{x})+g(\bm{x})\bm{u}\in\mathcal{T}_{\mathcal{C}}(\bm{x}),$ (2)
where $\mathcal{T}_{\mathcal{C}}(\bm{x})$ is the tangent cone to $\mathcal{C}$ in $\bm{x}$ [9, Definition 3.1].
While this characterization was groundbreaking, working with the tangent cone $\mathcal{T}_{\mathcal{C}}(\bm{x})$ can be cumbersome. Fortunately, when the set $\mathcal{C}$ satisfies certain regularity and transversality conditions (see [5] for details), the tangent cone on the boundary is simplified.
Property 1 (MFCQ [31])
Let $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ be continuously differentiable and denote its channels by $h_{i}:\mathcal{D}\to\mathbb{R}$. We say $h$ satisfies the MFCQ if
$\forall\bm{x}\in\mathcal{D},\quad\exists\bm{z}\in\mathbb{R}^{n},\quad\forall i\in\mathcal{I}(\bm{x}),\quad\frac{\partial h_{i}}{\partial\bm{x}}\bm{z}>0,$ (3)
where $\mathcal{I}(\bm{x})=\{i\mid h_{i}(\bm{x})=0\}$ is the set of active constraints.
Proposition 1 ([5, Proposition 4.3.7])
Let $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ be continuously differentiable. If $h$ satisfies the MFCQ, then
$\forall\bm{x}\in\partial\mathcal{C},\quad\mathcal{T}_{\mathcal{C}}(\bm{x})=\left\{\bm{z}\in\mathbb{R}^{n}\bigg|\min_{i\in\mathcal{I}(\bm{x})}\frac{\partial h_{i}}{\partial\bm{x}}\bm{z}\geq 0\right\},$ (4)
where $\mathcal{C}=\{\bm{x}\in\mathcal{D}\mid\min_{i\in\{1,\ldots,n_{c}\}}h_{i}(\bm{x})\geq 0\}$.
Theorem 1 and Proposition 1 imply the following result.
Theorem 2
Let $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ be a continuously differentiable function that satisfies the MFCQ. The set
$\mathcal{C}=\left\{\bm{x}\in\mathcal{D}\;\Big|\;\min_{i\in\{1,\ldots,n_{c}\}}h_{i}(\bm{x})\geq 0\right\},$ (5)
is control invariant if and only if $\forall\bm{x}\in\partial\mathcal{C}$, $\exists\bm{u}\in\mathcal{U}$, such that
$\min_{i\in\mathcal{I}(\bm{x})}\left[L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}\geq 0\right],$ (6)
where $L_{f}h_{i}$ and $L_{g}h_{i}$ are the Lie derivatives of $h_{i}$ along $f$ and $g$, respectively.
@@
II-B Control Barrier Functions
Control barrier functions certify the control invariance of their superlevel set. Traditionally, CBFs are real-valued functions [3] and, when multiple CBFs need to be imposed simultaneously, the control-sharing property presented in [44] ensures their compatibility. In this work, we combine these ingredients to present a general, vector-valued CBF.
Definition 2
@@ II-B Control Barrier Functions
A continuously differentiable function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ is a control barrier function (CBF) if there exists $\alpha\in\mathscr{K}$ for which, $\forall\bm{x}\in\mathcal{C}$, $\exists\bm{u}\in\mathcal{U}$, such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}+\alpha\big(h_{i}(\bm{x})\big)\right]\geq 0,$ (7)
where $\mathcal{C}$ is given in (5).
Remark 1
While we stated Definition 2 using a unique class $\mathscr{K}$ function $\alpha$ to simplify the presentation, it is equivalent to considering a different $\alpha_{i}\in\mathscr{K}$ for each $h_{i}$. Tuning each $\alpha_{i}$ allows the user to modulate the performance of the closed-loop system while retaining the safety guarantees as long as (7) is satisfied with $\alpha\to\alpha_{i}$.
@@
With this definition, the conditions of Theorem 2 can be extended to the entire set $\mathcal{C}$, as shown in the next result.
Theorem 3
[…]
  1. 1.
    If the function $h$ is a CBF, then $\mathcal{C}$ is control invariant.
  2. 2.
    Assume (a) $\mathcal{C}$ is compact and (b) there exists a continuous selection $\mu:\mathcal{C}\to\mathcal{U}$ such that $\mu(\bm{x})\in\mathcal{K}(\bm{x})$. If $\mathcal{C}$ is control invariant, then $h$ is a CBF.
Proof:
Although the proof is based on [3], they showed the result for real-valued barrier functions. Our result concerns vector-valued control barrier functions.
We begin by showing statement 1. Assume $h$ is a CBF and let $\alpha\in\mathscr{K}$ satisfy the CBF definition. Let $\bm{x}\in\partial\mathcal{C}$ be given and pick $\bm{u}\in\mathcal{U}$ such that (7) holds. Let $i\in\mathcal{I}(\bm{x})$ and note that $0\leq L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}+\alpha\big(h_{i}(\bm{x})\big)=L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}+\alpha(0)=L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\bm{u}$. We conclude by Theorem 2 that $\mathcal{C}$ is control invariant.
Next, we show statement 2. Let $\mathcal{C}$ be control invariant and pick $i\in\{1,\ldots,n_{c}\}$. Define the set-valued map
$\mathcal{C}_{i}(c)=\{\bm{x}\in\mathcal{C}\mid 0\leq h_{i}(\bm{x})\leq c\},$ (8)
and the value function $\hat{\alpha}_{i}:[0,\infty)\to\mathbb{R}$ by
$\hat{\alpha}_{i}(c)=-\inf_{\bm{x}\in\mathcal{C}_{i}(c)}\left[L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\mu(\bm{x})\right].$ (9)
For any $c\geq 0$, the set $\mathcal{C}_{i}(c)\subset\mathcal{C}$ is compact. Next, we argue the set-valued map $\mathcal{C}_{i}$ is upper and lower hemicontinuous at 0: Let $\{c_{k}\}\subset[0,\infty)$ be a sequence that converges to 0 and let $\{\bm{x}_{k}\}\subset\mathcal{D}$ be a sequence that converges to $\bm{x}\in\mathcal{D}$ satisfying $\bm{x}_{k}\in\mathcal{C}_{i}(c_{k})$. Noting that $0\leq h_{i}(\bm{x}_{k})\leq c_{k}$ for all $k\in\mathbb{N}$, it follows by continuity of $h_{i}$ that $0\leq h_{i}(\bm{x})\leq 0$, and thus $\bm{x}\in\mathcal{C}_{i}(0)$, concluding that $\mathcal{C}_{i}$ is upper hemicontinuous at 0; For lower hemicontinuity, pick any $\bm{x}\in\mathcal{C}_{i}(0)$ and any sequence $\{c_{k}\}\subset[0,\infty)$ that converges to 0. Clearly, $\bm{x}\in\mathcal{C}_{i}(c_{k})$ for all $k\in\mathbb{N}$, so we can create the trivial sequence $\{\bm{x}_{k}\}\subset\mathcal{D}$ with $\bm{x}_{k}=\bm{x}$ to conclude that $\mathcal{C}_{i}$ is lower hemicontinuous at 0. With this, it follows by Berge’s maximum theorem [8] that the value function $\hat{\alpha}_{i}$ is continuous at 0. By construction, $\hat{\alpha}_{i}$ is non-decreasing. Next, let $\bm{x}\in\mathcal{C}_{i}(0)\subset\mathcal{C}$ be given and note that since $\mu(\bm{x})\in\mathcal{K}(\bm{x})$, $L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\mu(\bm{x})\geq 0$, and it follows that $\hat{\alpha}_{i}(0)\leq 0$. There always exists $\alpha_{i}\in\mathscr{K}$ that upper-bounds $\hat{\alpha}_{i}$. Since $i\in\{1,\ldots,n_{c}\}$ was arbitrary, let us define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}$. Finally, let $\bm{x}\in\mathcal{C}$ be given and recall $\mu(\bm{x})\in\mathcal{U}$. For any $i\in\{1,\ldots,n_{c}\}$,
$\dot{h}_{i}\big(\bm{x},\mu(\bm{x})\big)$ $=L_{f}h_{i}(\bm{x})+L_{g}h_{i}(\bm{x})\mu(\bm{x})$
$\geq\inf_{\bm{x}_{0}\in\mathcal{C}_{i}\big(h_{i}(\bm{x})\big)}\left[L_{f}h_{i}(\bm{x}_{0})+L_{g}h_{i}(\bm{x}_{0})\mu(\bm{x}_{0})\right]$
$\geq-\hat{\alpha}_{i}\big(h_{i}(\bm{x})\big)$
$\geq-\alpha_{i}\big(h_{i}(\bm{x})\big)$
$\geq-\alpha\big(h_{i}(\bm{x})\big).$
We conclude that $h$ is a CBF. ∎
Remark 2
When $\mathcal{C}$ is control invariant, the set $\mathcal{K}(\bm{x})\neq\emptyset$ for every $\bm{x}\in\mathcal{C}$. Therefore there always exists a selection $\mu:\mathcal{C}\to\mathcal{U}$ such that $\mu(\bm{x})\in\mathcal{K}(\bm{x})$. Statement 2 of Theorem 3 requires at least one such selection be continuous. Moreover, in the special case that $n_{c}=1$, the following statement also holds: Assume (a) $\mathcal{C}$ is compact and (b) $\mathcal{U}$ is compact. If $\mathcal{C}$ is control invariant, then $h$ is a CBF.
@@
CBFs can be used to design add-on modules that enforce constraint satisfaction by filtering the control input. Specifically, let $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ be a CBF such that $\mathcal{C}\subset\mathcal{X}$ and consider the closed-loop system
$\dot{\bm{x}}=f(\bm{x})+g(\bm{x})\beta(\bm{x}),$ (10)
@@
$\beta(\bm{x})=\underset{\bm{u}\in\mathcal{K}_{\alpha}(\bm{x})}{\operatorname{argmin}}\>\|\bm{u}-\kappa(\bm{x})\|^{2},$ (11)